How much energy can be stored in a capacitor with two parallel plates, each with an area of 64.0 cm2 and separated by a gap of 1.30 mm, filled with porcelain whose dielectric constant is 7.0, and holding equal and opposite charges of magnitude 420. μC?


1
Expert's answer
2021-04-29T10:42:33-0400

Find the capacitance:


"C=\frac{k\epsilon_0A}{d}."

The energy stored is


"E=\frac{Q^2}{2C}=\frac{Q^2d}{2k\epsilon_0A}=28.9\text{ kJ}."